# Transfer and routing of Gaussian states through quantum complex networks with and without community structure

### v1

@@ IV.1 Role of community structure
As an example of a real complex network with community structure we studied the routing properties of Zachary’s Karate club network [38]. A network of QHOs with the unweighted karate club structure was constructed with the same parameters as in Sec. III.2, and transfer of a squeezed vacuum $\left|r=1,\varphi=0\right>$ was again considered. Figure 8 shows the mean transfer fidelities within and between the communities for the five slowest normal modes. In contrast to the random modular networks in Sec. III.2, which had a more pronounced community structure, we find that while communication between the two communities is only feasible over the two slowest normal modes, both communities have additional normal modes over which at least some communication is possible. For even higher normal modes there are in general only a few or no pairs which reach any reasonable fidelity.
@@
Figure 11: Fraction of pairs for which $F>0.8$ for all normal modes of the networks presented in the article. The networks have the same parameters as in sec. III.2. For the random networks only a single realization is presented here. We see that the modular networks have a handful of decent normal modes, while the adjective-noun and ER networks mainly allow transfer over the center of mass mode.
@@
Fig. 11 highlights another feature of networks that is important for routing: the adjective-noun network has a very dense frequency distribution, while the modular networks have some separation between the suitable normal modes. Although the spectrum of the adjective-noun network has large gaps at the highest frequencies, only few nodes couple to those frequencies, limiting their usefulness for routing.
IV.2 Comparison with randomized counterparts
@@
The first quantity attempts to capture the importance of a network node as a point of contact. Let $F_{i,j}^{n}$ be the transfer fidelity when external users couple to network nodes $i$ and $j$ and transfer happens through the $n$-th normal mode and let $t^{n}$ be the associated transfer time, inversely proportional to the normal mode frequency but independent of $i$ and $j$. Consider the quantity $C_{i,j}:=\sum_{n=1}^{N}F_{i,j}^{n}/t^{n}$, where $N$ is the network size. Notice that $C_{i,j}=C_{j,i}$. To focus on the importance of a single network node, we consider the quantity
@@
The value of $C_{i}$ for some node $i$ is relatively high if it can support higher fidelity transfers at faster speeds compared to the other nodes.
@@
Results for the two networks are shown in Figs. 12 and 13. Like previously a low degree is found to be detrimental, however we also observe a new behavior where nodes with a particularly high degree reach lower node capacities than nodes with an intermediate degree. Overall the behavior can be expected to be robust to changes in network structure as it is observed in two quite different networks, although the effect is particularly strong in the adjective-noun network. Considering results of Sec. III.1, the nodes with the highest degree can be expected to support high fidelity transfer mostly with the slowest normal mode while nodes with intermediate degree might have a more even spread of transfer fidelities across the available normal modes.
@@
While the best pairs are not necessarily free of overlapping users, especially for larger networks we expect $C_{max}$ to be a good approximation of the optimal set of transfers that can happen simultaneously. We compare this quantity to that achieved with randomized counterparts for both networks. The random counterparts preserve some but not all features of the original networks as explained in Appendix D.
@@ V Conclusions
We leave open the question of which other features of the network may be relevant to routing, and how properties such as network diameter, degree distribution or clustering affect the transfer performance. In particular, if one can construct an arbitrary network, how should it be designed to maximize routing performance? When is a network suitable only for sequential high fidelity transfer better than a network that can support relatively high fidelity transfer over multiple normal modes, and vice versa? To systematically explore this and the impact of the previously mentioned network properties calls for further work on quantifying the capacity of a network for both state transfer and for routing. Ideally, it should have a clear physically motivated definition, produce a single real number—to facilitate ranking the networks—with clear interpretation and be applicable to a variety of physical systems beyond just continuous variable Gaussian states. It could be expected to single out networks with many useful normal modes. The tentative quantities we have proposed in Sec. IV.2 fulfill some of these requirements already, however especially in the case of Gaussian sates there seems to be no obvious way to make the capacity state independent. Furthermore, the ability to trade speed for improved fidelity should be taken into account in some appropriate manner. We hope to tackle this challenge in the future.
@@ Appendix A Transferred states
where $y=\cosh{(2r)}$, $z_{\text{cos}}=\cos{(\varphi)}\sinh{(2r)}$ and $z_{\text{sin}}=-\sin{(\varphi)}\sinh{(2r)}$ and where $\Re(\alpha)$ and $\Im(\alpha)$ are the real and imaginary parts of the displacement, respectively. For squeezed states $\left|r,\varphi\right>$ we set $\alpha=0$ and for coherent states $\left|\alpha\right>$ we set $r=0$.
@@ Appendix B State transfer in short oscillator chains with linear couplings
As pointed out in Sec. II.2, the cases relevant for our purposes concern permutations in a chain of three oscillators. Without a loss of generality we may assume that the normal mode frequency is $\omega_{0}$ and that the sender and the receiver have tuned their frequencies to match it. Focusing first on applying twice the state swap between two oscillators, $\mathbf{P}=\pm\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, $\mathbf{K}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$ and $\mathbf{\Omega}=\{\sqrt{\omega_{0}^{2}-g},\sqrt{\omega_{0}^{2}+g}\}$. The freedom to use either sign for the matrix $\mathbf{P}$ is because it will be used twice, and the second application will restore the original phase if necessary. The diagonalization is valid when $g<\omega_{0}^{2}$. From Eq. (8) we get $\cos(t\sqrt{\omega_{0}^{2}-g})=-\cos(t\sqrt{\omega_{0}^{2}+g})=1$ for positive $\mathbf{P}$ and $-1$ otherwise whereas in both cases $\sin(t\sqrt{\omega_{0}^{2}-g})=\sin(t\sqrt{\omega_{0}^{2}+g})=0$. Taking $\omega_{0}$ to be fixed, we solve for $g$ and $t$ and find a two-parameter family of solutions $g=\frac{(1+c^{\prime})^{2}-c^{\prime\prime 2}}{(1+c^{\prime})^{2}+c^{\prime\prime 2}}\omega_{0}^{2}$, $t=\sqrt{\frac{1}{2}+c^{\prime}(1+c^{\prime}/2)+c^{\prime\prime 2}/2}\pi\omega_{0}^{-1}$ where $c^{\prime}\geq c^{\prime\prime}\geq 1$ are integers. Transfer time $t$ increases with $c^{\prime}$ and $c^{\prime\prime}$ whereas setting $c^{\prime\prime}=ac^{\prime}+b$ gives $\lim_{c^{\prime}\rightarrow\infty}g=\frac{1-a^{2}}{1+a^{2}}$. We are interested in particular in the solutions where $g$ tends to 00 as transfer time increases as this reduces the detrimental effect of the other normal modes, which corresponds to choosing $c^{\prime}=c^{\prime\prime}:=c_{2}$. Therefore we will use
@@ Appendix C Asymptotic behavior of transfer times and coupling strengths
Appendix D Network randomization algorithms
Complex networks can exhibit a wealth of non-trivial topological features, from heterogeneous degree distributions to degree-degree correlations and clustering. In order to assess the impact of these properties on transport efficiency, we constructed, for both real network considered, several ensembles containing 100 networks each in which some of these properties are preserved. The randomisations were carried out using the code developed for Ref. [50] (repository available at [51]). For any input network, the program generates a maximally random network with the same degree sequence. Other properties of the input graph can be preserved on demand as well (for instance, degree correlations, embodied in the joint degree distribution $P(k,k^{\prime})$, and/or the clustering spectrum $c(k)$). The four 100-network ensembles were generated by preserving
@@ Appendix D Network randomization algorithms
    degree distribution $P(k)$ and clustering spectrum $c(k)$,
@@
    degree distribution $P(k)$, degree-degree correlations $P(k,k^{\prime})$, and clustering spectrum $c(k)$.

### v2

@@ 4.1 Role of community structure
As an example of a real complex network with community structure we studied the routing properties of Zachary’s Karate club network [41], a paradigmatic network with community structure. The network with its division to communities is shown in Fig. 9. A network of QHOs with the unweighted karate club structure was constructed with the same parameters as in Sec. 3.2, and transfer of a squeezed vacuum $\left|r=1,\varphi=0\right>$ was again considered. Figure 8 shows the mean transfer fidelities within and between the communities for the five slowest normal modes. In contrast to the random modular networks in Sec. 3.2, which had a more pronounced community structure, we find that while communication between the two communities is only feasible over the two slowest normal modes, both communities have additional normal modes over which at least some communication is possible. For even higher normal modes there are in general only a few or no pairs which reach any reasonable fidelity.
@@
[…]
[trim=0cm 0cm 0.0cm 1cm,clip=true,width=0.48]figures/good_transfers_freq.pdf
Figure 11: Fraction of pairs for which $F>0.8$ for all normal modes of the networks presented in the article. The networks have the same parameters as in Sec. 3.2. For the random networks only a single realization is presented here. We see that the modular networks have a handful of decent normal modes, while the adjective-noun and ER networks mainly allow transfer over the center of mass mode.
@@
Figure 11 highlights another feature of networks that is important for routing: the adjective-noun network has a very dense frequency distribution, while the modular networks have some separation between the suitable normal modes. Although the spectrum of the adjective-noun network has large gaps at the highest frequencies, only few nodes couple to those frequencies, limiting their usefulness for routing.
4.2 Comparison with randomized counterparts
@@
The first quantity attempts to capture the importance of a network node as a point of contact. Let $F_{i,j}^{n}$ be the transfer fidelity when external users couple to network nodes $i$ and $j$ and transfer happens through the $n$-th normal mode and let $t^{n}$ be the associated transfer time, inversely proportional to the normal mode frequency but independent of $i$ and $j$. Consider the quantity $C_{i,j}:=\sum_{n=1}^{N}F_{i,j}^{n}/t^{n}$, where $N$ is the network size. Notice that $C_{i,j}=C_{j,i}$. To focus on the importance of a single network node, we consider node capacity $C_{i}$, defined as
@@
The value of $C_{i}$ for some node $i$ is relatively high if it can support higher fidelity transfers at faster speeds compared to the other nodes. To compare different networks’ capacities, these values should be then normalized. Here we average the capacities by the network size, $C_{i}/N$, however one could also consider averaging both $C_{i,j}$ and $C_{i}$ so that the overall result would be $C_{i}/N^{2}$.
Results for the two networks are shown in Figs. 12 and 13. Like previously a low degree is found to be detrimental, however we also observe a new behavior where nodes with a particularly high degree reach lower node capacities than nodes with an intermediate degree. Overall the behavior can be expected to be robust to changes in network structure as it is observed in two quite different networks, although the effect is particularly strong in the adjective-noun network. Considering results of Sec. 3.1, the nodes with the highest degree can be expected to support high fidelity transfer mostly with the slowest normal mode while nodes with intermediate degree might have a more even spread of transfer fidelities across the available normal modes.
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/nodecapacityKarate.pdf
@@
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/nodecapacityAdjNoun.pdf
@@
While the best pairs are not necessarily free of overlapping users, especially for larger networks we expect $C_{max}$ to be a good approximation of the optimal set of transfers that can happen simultaneously. We compare this quantity to that achieved with randomized counterparts for both networks. The random counterparts preserve some but not all features of the original networks as explained in App. E.
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.45]figures/maxcapacityKarate.pdf
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.45]figures/maxcapacityAdjnoun.pdf
@@ 5 Conclusions
We leave open the question of which other features of the network may be relevant to routing, and how properties such as network diameter, degree distribution or clustering affect the transfer performance. In particular, if one can construct an arbitrary network, how should it be designed to maximize routing performance? When is a network suitable only for sequential high fidelity transfer better than a network that can support relatively high fidelity transfer over multiple normal modes, and vice versa? To systematically explore this and the impact of the previously mentioned network properties calls for further work on quantifying the capacity of a network for both state transfer and for routing. Ideally, it should have a clear physically motivated definition, produce a single real number—to facilitate ranking the networks—with clear interpretation and be applicable to a variety of physical systems beyond just continuous variable Gaussian states. It could be expected to single out networks with many useful normal modes. The tentative quantities we have proposed in Sec. 4.2 fulfill some of these requirements already, however especially in the case of Gaussian states there seems to be no obvious way to make the capacity state independent. Furthermore, the ability to trade speed for improved fidelity should be taken into account in some appropriate manner. We hope to tackle this challenge in the future.
The modular structure of our model makes state transfer robust both to static disorder in the frequency of a node and to loss of random links [52]. What may cause poor transfer fidelity is weak coupling between the chosen normal mode and the sender or receiver. Here the prospects of noise having a positive role can be considered. This is a well-known phenomenon in excitation transport where it is known as noise-assisted transport [53, 54, 55] or environment-assisted transport [56, 57], but there are sporadic results concerning quantum state transfer as well [58, 59]. In the case at hand, balancing a weak coupling to a normal mode leads to a strong coupling in the network basis which can cause the off-resonant normal modes to disturb the transfer. Since it is known that time dependent modulation of the couplings to the network can realize a spectral filter blocking exchange of quantum information to certain frequencies [60], it is conceivable that judiciously engineered dynamical noise—that is to say, random time dependent variation with a specific distribution—in these couplings could improve transfer fidelity in such situations. A different approach would be to consider an open network, which would introduce environment-induced decoherence. Since its effect on transfer fidelity hasn’t received much attention, investigating it would be an interesting avenue of further research.
@@ Appendix A Transferred states
where $y=\cosh{(2r)}$, $z_{\text{cos}}=\cos{(\varphi)}\sinh{(2r)}$ and $z_{\text{sin}}=-\sin{(\varphi)}\sinh{(2r)}$, $n_{\text{th}}$ is the mean number of thermal excitations and $\Re(\alpha)$ and $\Im(\alpha)$ are the real and imaginary parts of the displacement, respectively. For squeezed states $\left|r,\varphi\right>$ we set $\alpha=0$ and for coherent states $\left|\alpha\right>$ we set $r=0$.
@@
For the non-Gaussian number states we cannot employ the description in terms of $\sigma$ and $\langle\mathbf{x}\rangle$. Instead they are presented in the discrete Fock basis, where each oscillator has a $d$-dimensional Hilbert space with basis vectors $\{\left|n\right>\}_{n=0}^{d}$. In theory $d=\infty$, but due to computational limitations $d<\infty$. The number states correspond to the basis vectors, so that an $n$-photon state is $\left|n\right>$.
@@ Appendix B State transfer in short oscillator chains with linear couplings
As pointed out in Sec. 2.2, the cases relevant for our purposes concern permutations in a chain of three oscillators. Without a loss of generality we may assume that the normal mode frequency is $\omega_{0}$ and that the sender and the receiver have tuned their frequencies to match it. Focusing first on applying twice the state swap between two oscillators, $\mathbf{P}=\pm\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, $\mathbf{K}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$ and $\mathbf{\Omega}=\{\sqrt{\omega_{0}^{2}-g},\sqrt{\omega_{0}^{2}+g}\}$. The freedom to use either sign for the matrix $\mathbf{P}$ is because it will be used twice, and the second application will restore the original phase if necessary. The diagonalization is valid when $g<\omega_{0}^{2}$. From Eq. (9) we get $\cos(t\sqrt{\omega_{0}^{2}-g})=-\cos(t\sqrt{\omega_{0}^{2}+g})=1$ for positive $\mathbf{P}$ and $-1$ otherwise whereas in both cases $\sin(t\sqrt{\omega_{0}^{2}-g})=\sin(t\sqrt{\omega_{0}^{2}+g})=0$. Taking $\omega_{0}$ to be fixed, we solve for $g$ and $t$ and find a two-parameter family of solutions $g=\frac{(1+c^{\prime})^{2}-c^{\prime\prime 2}}{(1+c^{\prime})^{2}+c^{\prime\prime 2}}\omega_{0}^{2}$, $t=\sqrt{\frac{1}{2}+c^{\prime}(1+c^{\prime}/2)+c^{\prime\prime 2}/2}\pi\omega_{0}^{-1}$ where $c^{\prime}\geq c^{\prime\prime}\geq 1$ are integers. Transfer time $t$ increases with $c^{\prime}$ and $c^{\prime\prime}$ whereas setting $c^{\prime\prime}=ac^{\prime}+b$ gives $\lim_{c^{\prime}\rightarrow\infty}g=\frac{1-a^{2}}{1+a^{2}}$. We are interested in particular in the solutions where $g$ tends to 00 as transfer time increases as this reduces the detrimental effect of the other normal modes, which corresponds to choosing $c^{\prime}=c^{\prime\prime}:=c_{2}$. Therefore we will use
@@ Appendix C Asymptotic behavior of transfer times and coupling strengths
Appendix D Details about numerical simulations
We begin by describing the case of Gaussian states, for which simulations are based on exact diagonalization and numerical real matrices. We construct the initial covariance matrix of the full system, diagonalize the total Hamiltonian to find the symplectic matrix that propagates it to the final covariance matrix, and pick the elements corresponding to the receiver. The vector of first moments is handled similarly as seen below. In all cases, both fidelity of Eq. (2) and efficiency of Eq. (3) can be calculated directly from the covariance matrices and first moments vectors of the sender and the receiver.
In more detail, the initial covariance matrices are based on Eqs. (6) and (7). The vacuum is recovered as a special case of the former by setting $n_{\text{th}}=0$ and $r=0$. The initial covariance matrix of the network is a product state of such vacua, except in Sec. 4.2 where it is the ground state of the network Hamiltonian (see, e.g., Equation (3.18) of [62]). There are no initial correlations between the sender, the receiver and the network. Consequently the full initial covariance matrix is formed simply by a direct sum of the three initial covariance matrices, followed by a reordering according to the chosen order of operators.
Let $\sigma(\mathbf{x}(0))$ be this initial numerical covariance matrix where $\mathbf{x}(0)$ is the column vector of operators at time $t=0$. Then $\sigma(\mathbf{x}(t))=\mathbf{S}(t)\sigma(\mathbf{x}(0))(\mathbf{S}(t))^{\top}$. For first moments vector $\langle\mathbf{x}(t)\rangle=\mathbf{S}(t)\langle\mathbf{x}(0)\rangle$. The construction of the symplectic matrix $\mathbf{S}(t)$ is essentially described in the first paragraph of App. B. Specifically, we start from the matrix of Eq. (1), add the terms for the sender and receiver frequencies with a direct sum, and include the linear coupling terms. The resulting matrix is real and symmetric and can be diagonalized by numerically finding its eigendecomposition which then determines $\mathbf{S}(t)$ through Eq. (8). The elements of the final covariance matrix and first moments vectors for the receiver are picked from $\sigma(\mathbf{x}(t))$ and $\langle\mathbf{x}(t)\rangle$, respectively.
In Sec. 2.1 we also consider the transfer of a number state in the ideal case as an example of a non-Gaussian state transfer. The procedure for their simulation is substantially different from the Gaussian case, utilizing a finite Hilbert space whose dimension is truncated at a suitable number. For our simulations we chose $d=20$: the states considered go up to $\left|n=15\right>$, and even for the highest photon initial states we found that the probabilities of states with $n>20$ in the received state were insignificant, below $10^{-6}$.
[…]
To calculate the fidelity in the single mode Gaussian case we use [64]
$\mathcal{F}(\rho_{1},\rho_{2})=$ $\frac{1}{\sqrt{\Delta+\Lambda}-\sqrt{\Lambda}}\times$ (14)
$e^{-\frac{1}{2}(\langle\mathbf{x}_{2}\rangle-\langle\mathbf{x}_{1}\rangle)^{T}(\sigma_{1}+\sigma_{2})^{-1}(\langle\mathbf{x}_{2}\rangle-\langle\mathbf{x}_{1}\rangle)},$
[…]
Efficiency requires the final value of various quantities. Displacement $\alpha$ and squeezing parameter $r$ can be solved from the final covariance matrix of the receiver using Eqs. (6). For the number states the expectation value for the number of excitations $\langle n\rangle$ is calculated from the density matrix of the receiver $\rho_{R}$ with $\langle n\rangle=\mathrm{Tr}(n\rho_{R})$, with $n=a^{\dagger}a$ being the number operator. Logarithmic negativity is defined to be
$\mathrm{E}(\sigma)=\max\{0,-\log_{2}{2\tilde{d}_{-}}\}.$ (15)
where $\tilde{d}_{-}$ is the smaller of the two so-called symplectic eigenvalues of a two-mode covariance matrix $\sigma$. Its value is calculated from the elements of $\sigma$ using Eq. (13) of [66].
Appendix E Network randomization algorithms
Complex networks can exhibit a wealth of non-trivial topological features, from heterogeneous degree distributions to degree-degree correlations and clustering. In order to assess the impact of these properties on transport efficiency, we constructed, for both real network considered, several ensembles containing 100 networks each in which some of these properties are preserved. The randomisations were carried out using the code developed for Ref. [67] (repository available at [68]).
For any input network, the program generates an ensemble of 100 random networks preserving either
@@ Appendix E Network randomization algorithms
    degree distribution $P(k)$ and clustering spectrum $c(k)$ or
@@
    degree distribution $P(k)$, degree-degree correlations $P(k,k^{\prime})$, and clustering spectrum $c(k)$
of the original network. Here, degree distribution $P(k)$ is the probability distribution of a node having degree $k$, degree-degree correlations (joint degree distribution) $P(k,k^{\prime})$ are the probabilities of nodes with degrees $k$ and $k^{\prime}$ having a link between them, and clustering spectrum $c(k)$ is the spectrum of mean local clustering coefficient of nodes of a given degree $k$. Local clustering coefficient of a node is the probability that a pair of its neighbours are also neighbours of each other [69].